Two hundred recent college graduates were surveyed about their social media connections. Of those surveyed, 120 used Facebook, 80 used LinkedIn, and 90 used Instagram. Moreover, 55 used both Facebook and Instagram, 45 used both LinkedIn and Facebook, 45 used both Instagram and LinkedIn, and 30 used all three platforms.
How many of the individuals surveyed used Instagram, but did not use either Facebook or LinkedIn?

Answer by ikleyn(52781)

.

From n(Instagram), we should subtract n(Instagram AND Facebook) and n(Instagram AND LinkedIn) and add n(all three)


    answer = n(Instagram) - n(Instagram AND Facebook) - n(Instagram AND LinkedIn) + n(all three) = 90 - 55 - 45 + 30 = 20.

Solved.